OpenStudy (anonymous):

find the value of a so that the lines (a+b)x + (a-b)y =4 and x+y = 1 are the same lines

hero (hero):

@cgoel, So we know x + y = 1 is a given line right?

OpenStudy (anonymous):

yes

hero (hero):

And we want (a + b)x + (a - b)y = 4 to be the same line right?

OpenStudy (anonymous):

correct

hero (hero):

So what we can do is multiply both sides of x + y = 1 by four. And if we do that we get 4x + 4y = 4. Do you agree?

OpenStudy (anonymous):

yes i agree, that makes sense

hero (hero):

So if both lines are the same, and both lines equal the same value, then the coefficients of each term must also be the same as well right?

hero (hero):

In other words a + b = 4 a - b = 4 Do you agree?

OpenStudy (anonymous):

sure, definatley

hero (hero):

Okay, and we only need to solve for variable a at this point. Do you see a way to do that?

OpenStudy (anonymous):

um, no...

hero (hero):

If we added both equations together, we could eliminate variable b and solve for a. It works like this. We simply add the corresponding like terms of each equation together like so: (a + a) + (b - b) = 4 + 4

hero (hero):

So we end up with 2a = 8

hero (hero):

Do you see that?

OpenStudy (anonymous):

oh yeah, now i see, so a =4. great! i also have another question

hero (hero):

Just out of curiosity, do you know what b equals?